Actions
2017-02-28 physics-shifter¶
Tasks:- Before Thursday - recovery from quench
- Wednesday-Thursday 23rd Feb - SSU and SSD only, powered in flip mode; momentum scan for magnet alignment
- Thursday 24th Feb - Sunday 26th Feb - magnets ramped to 2016/05 1
- momentum scan then emittance scan
- diffuser 4 stuck on highest emittance setting
- Sunday 26th Feb - Tuesday 28th Feb - ISIS down for failure of dipole 9
- Emittance Reduction
Magnet alignment¶
- Noted issues with track extrapolation from detectors
- For SSs only data, SSD was marked as solenoid mode in the reconstruction software; resolved, awaiting reconstruction reprocessing
- For 2016/05 1 data, SSD Trim coils were marked as 0 current; resolved, awaiting reconstruction reprocessing
Emittance Reduction¶
- Note that DiffMask8-3-140+M3-TSpecial Run is misaligned
Tag | Runs | lmc1234 sum | TOF1 Triggers | TOF2 Triggers | time taken | number in muon peak | number at tof2 | number passing all cuts | number passing all cuts / p in tku [MeV/c^-1] | number passing all cuts / tof01 [ns^-1] |
DiffMask8-3-140+M3-TSpecial Run | [9209] | 610282 | 150483 | 46090 | 1 hrs 58 mins | 49146 | 18522 | 13312 | 1438.62 | 25815.0 |
10-140+M3-Test3 | [9203] | 326423 | 245332 | 37004 | 0 hrs 53 mins | 64362 | 12404 | 9521 | 1099.26 | 20965.0 |
Sample output¶
Updated by Rogers, Chris about 6 years ago ยท 5 revisions